OpenStudy (ineedhelp10):

for a line with the points (3,3) and (4,k) what are the values of k that would make the slope undefined

OpenStudy (anonymous):

Do you know what makes a slope undefined?

OpenStudy (ineedhelp10):

yeah its when the slope ends up being error

OpenStudy (anonymous):

It's when the denominator is equal to 0. That is, there is no change in x. A vertical line

OpenStudy (anonymous):

do you know the equation for the slope of a line given two points?

OpenStudy (ineedhelp10):

yeah its y2-y1/x2-x1

OpenStudy (anonymous):

ok so plug in the numbers you have into that equation

OpenStudy (anonymous):

There is a change in x of 1 unit

OpenStudy (anonymous):

so there is no y value that can make it undefined

OpenStudy (ineedhelp10):

okay but what do i set it equal to?

OpenStudy (anonymous):

(k-3)/(4-3) = (k-3)/1

OpenStudy (anonymous):

No k value makes the slope undefined

OpenStudy (anonymous):

Do you see what I mean?

OpenStudy (ineedhelp10):

would it be k-3?

OpenStudy (anonymous):

the slope is undefined if x2-x1 = 0, correct?

OpenStudy (ineedhelp10):

wait would it be all real numbers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

for the slope to be undefined, x2-x1 must equal 0, right?

OpenStudy (anonymous):

(y2-y1)/(x2-x1) (x2-x1) is the denominator. If (x2-x1) = 0, then the slope is undefined.

OpenStudy (anonymous):

right?

OpenStudy (ineedhelp10):

okay so then what?

OpenStudy (anonymous):

well x2=4 and x1=3 so (x2-x1) = 1 No y value will change that. So the slope cannot be undefined

OpenStudy (ineedhelp10):

okay so then what would that make the value of k?

OpenStudy (anonymous):

Are you sure you typed the problem right? It would make more sense if it asked about (3,3) and (k.4)

OpenStudy (ineedhelp10):

yeah i typed it right

OpenStudy (anonymous):

changing k won't change the value of the denominator. So there is no value of k that would make the slope undefined.

OpenStudy (ineedhelp10):

okay you're so sure

OpenStudy (anonymous):

it's either a trick question or your teacher made a mistake

OpenStudy (ineedhelp10):

and for the slope to be 0 would it have to be 3

OpenStudy (anonymous):

correct

OpenStudy (ineedhelp10):

okay thank you so much!!

OpenStudy (anonymous):

No problem

OpenStudy (ineedhelp10):

wait for the undefined would it be all real numbers?

OpenStudy (anonymous):

That would mean that the slope is always undefined It is never undefined

OpenStudy (ineedhelp10):

so then what can i say as my final answer?

OpenStudy (anonymous):

I would say that no value of k makes the slope undefined

OpenStudy (ineedhelp10):

okay thehn thank you again!!

OpenStudy (anonymous):

Of course

OpenStudy (ineedhelp10):

wait what can i say for my explanation lol my bad

OpenStudy (ineedhelp10):

do you know?

OpenStudy (anonymous):

There is no value of k that makes the slope undefined.